# Beyond Chunking: Discourse-Aware Hierarchical Retrieval for Long Document Question Answering

### v3

@@ 4.3 Discussions
We evaluate our method on NarrativeQA, which contains exceptionally long documents (average: 51,372 words, maximum: 346,902 words) that exceed most generative models’ context limits. DISRetrieval achieves superior performance across all metrics, outperforming the widely-used flatten-chunk baseline by +1.35% BLEU, +1.84% ROUGE, and +1.51% METEOR. Notably, both DISRetrieval and RAPTOR substantially outperform flatten-based methods, confirming that structured approaches are more effective than flat for extremely long document retrieval. However, DISRetrieval’s discourse-aware structure provides consistent advantages over RAPTOR’s semantic clustering approach.
@@
To evaluate the importance of precise evidence retrieval, we compare three context settings using Llama3.1-8B-Instruct: (i) golden evidence, (ii) full document, and (iii) full document augmented with retrieved evidence (highlighted with [EVIDENCE] … [/EVIDENCE] markers). Golden evidence achieves the highest F1 score of 50.71% with only 129 words on average, significantly outperforming the full document approach (48.81% with 4,170 words). This demonstrates that concise yet accurate retrieval is more effective than using substantially larger context. Furthermore, augmenting the full document with our retrieved evidence yields consistent improvements of 0.75-1.02% across all settings, confirming that precise retrieval enhances performance even when combined with complete document access. These results underscore that effective long-document QA systems should prioritize retrieval quality over context quantity.
@@
We investigate whether model scale affects the quality of discourse-aware tree construction by comparing different LLMs. As shown in Table 4, smaller models like Llama-3.1-8B, Qwen2.5-7B, and Mistral-7B achieve comparable performance to larger models, with differences under 0.5%. Notably, Llama-3.1-8B achieves the best recall at 400-word context length. These results indicate that the effectiveness of our discourse-aware retrieval does not heavily depend on LLM scale. This finding is particularly valuable as it enables flexible model selection without requiring extremely large models, making our approach more practical and cost-effective.
@@
Llama-3.1-8B 28.66 61.76 24.74 68.30 21.68 72.92
Qwen2.5-7B 28.32 61.09 24.75 68.11 21.69 72.53
Mistral-7B 28.17 60.86 24.65 68.41 21.65 72.86
GPT-4o-mini 28.27 61.13 24.66 68.31 21.68 72.84
Deepseek-v3 28.89 61.91 25.02 68.57 21.82 72.91
@@
We examine how discourse parser capabilities affect downstream performance by training parsers on varying amounts of data (0-100%). As shown in Figure 4, both retrieval recall and answer F1 scores improve consistently as parser training data increases across all context lengths. These results confirm that parser quality directly impacts the effectiveness of our discourse-aware tree structure. This suggests that improving discourse parsing capabilities is a promising direction for enhancing our approach’s performance.
@@ 5 Conclusion
In this paper, we presented DISRetrieval, a discourse-aware hierarchical retrieval framework that systematically incorporates rhetorical structure theory into long document question answering. Through three key innovations—sentence-level discourse parsing adapted for lengthy texts, LLM-enhanced node representations, and structure-aware evidence selection—our approach demonstrates substantial improvements over existing methods across multiple challenging benchmarks. Comprehensive experiments on QASPER, QuALITY, and NarrativeQA demonstrate consistent performance improvements across varying context lengths, embedding models, and generation architectures. Our ablation studies confirm that discourse structure provides more principled document organization than semantic clustering approaches, with the framework naturally adapting to different document types through multi-granularity retrieval. Beyond performance improvements, this work demonstrates that linguistically-grounded approaches remain valuable in the era of LLMs, offering principled alternatives to purely data-driven methods. Our framework opens new directions for incorporating structured linguistic knowledge into neural retrieval systems, with potential applications in legal document analysis, scientific literature review, and educational content processing.
@@
While DISRetrieval demonstrates substantial improvements across multiple datasets and architectures, we acknowledge several limitations that point to promising future directions. First, our discourse parser’s performance inherently bounds the overall system effectiveness, as shown in Figure 4 where parser quality directly impacts downstream performance. However, our results demonstrate that even with a straightforward EDU-to-sentence adaptation, significant gains are achieved, validating the core hypothesis that discourse structure benefits long document retrieval. Future work will explore more sophisticated discourse parsing techniques, including domain-adaptive parsers and multi-granularity discourse analysis. Second, our adaptive summarization strategy using threshold $\tau$ is relatively simple but proves effective across diverse document types ($\tau$=0 for academic papers, $\tau$=50 for narratives). While this approach successfully balances computational efficiency with representation quality, more sophisticated dynamic thresholding based on content complexity and hierarchical position could further optimize the trade-off. Third, current evaluation metrics, though comprehensive across three challenging datasets, may not fully capture the nuanced benefits of discourse-aware retrieval, such as structural coherence preservation and hierarchical information flow. We plan to develop evaluation frameworks that better assess discourse-aware retrieval quality across multiple dimensions.
Ethical Statement
@@
This research adheres to ethical principles in natural language processing research and does not raise significant ethical concerns. Our work focuses on improving document retrieval and question answering systems through discourse-aware hierarchical modeling, which has potential positive societal impacts by enhancing information access and comprehension. The datasets used in our experiments (QASPER, QuALITY, and NarrativeQA) are publicly available benchmark datasets that have been previously vetted by the research community and do not contain sensitive personal information or harmful content. Our discourse parsing approach operates on linguistic structures rather than content semantics, minimizing risks of bias amplification or misuse. The proposed DISRetrieval framework is designed as a general-purpose retrieval enhancement technique that can benefit various applications requiring long document understanding, such as academic research assistance, educational content processing, and legal document analysis. We acknowledge that like any information retrieval system, our method could potentially be misused if applied to spread misinformation or manipulate access to information, but such concerns are not specific to our approach and apply broadly to information retrieval technologies. We encourage responsible deployment of our system with appropriate safeguards and human oversight, particularly in high-stakes applications. All experiments were conducted using publicly available computational resources and open-source tools, ensuring reproducibility and transparency in our research process.
@@ B.1 Dataset Specifications
Table 6: Detailed infomation of the datasets used in our experiments.
@@
The NarrativeQA dataset is a significant advancement in evaluating machine reading comprehension. Unlike traditional datasets that focus on short passages, NarrativeQA emphasizes understanding long-form narratives, such as entire books or full-length movie scripts. One of the defining features of NarrativeQA is the length of its documents. On average, each document contains 51,372 words. Notably, some individual documents extend up to 346,902 words, making it one of the longest reading comprehension datasets available. More detailed information is presented in Table 6.
B.2 Model Specifications
@@ B.3 Implementation Details
B.4 Experiments Compute Resources
@@ B.5 Prompts
Given text: {left child node text} {input cild node text}
@@
User: Using the folloing information: {context}. Answer the following question in less than 5-7 words, if possible. For yes or no question, only return ’yes’ or ’no’.

### v4

@@ 4.3 Discussions
We evaluate our method on NarrativeQA, which contains exceptionally long documents (average: 51,372 words, maximum: 346,902 words) that exceed most generative models’ context limits. DISRetrieval achieves superior performance across all metrics, outperforming the widely-used flatten-chunk baseline by +1.15% BLEU, +1.89% ROUGE, and +1.46% METEOR. Notably, both DISRetrieval and RAPTOR substantially outperform flatten-based methods, confirming that structured approaches are more effective than flat for extremely long document retrieval. However, DISRetrieval’s discourse-aware structure provides consistent advantages over RAPTOR’s semantic clustering approach.
@@
To evaluate the importance of precise evidence retrieval, we compare three context settings using Llama3.1-8B-Instruct: (i) golden evidence, (ii) full document, and (iii) full document augmented with retrieved evidence (highlighted with [EVIDENCE] … [/EVIDENCE] markers). Golden evidence achieves the highest F1 score of 50.71% with only 129 words on average, significantly outperforming the full document approach (48.81% with 4,170 words). This demonstrates that concise yet accurate retrieval is more effective than using substantially larger context. Furthermore, augmenting the full document with our retrieved evidence yields consistent improvements of 0.73-1.39% across all settings, confirming that precise retrieval enhances performance even when combined with complete document access. These results underscore that effective long-document QA systems should prioritize retrieval quality over context quantity.
@@
[…]
@@
Llama-3.1-8B 28.13 60.75 24.62 67.98 21.58 72.61
Qwen2.5-7B 28.54 61.08 24.73 68.13 21.76 72.71
Mistral-7B 28.51 61.57 24.68 68.53 21.54 72.50
GPT-4o-mini 28.29 61.63 24.65 68.42 21.49 72.65
Deepseek-v3 28.66 61.80 24.77 68.66 21.72 72.69
@@
We examine how discourse parser capabilities affect downstream performance by training parsers on varying amounts of data (0-100%). As shown in Figure 4, both retrieval recall and answer F1 scores improve consistently as parser training data increases across all context lengths. Notably, despite our parser being trained only on RST-DT (news-text-dominant), DISRetrieval consistently outperforms all baselines across diverse genres (research papers, fiction, movie scripts), demonstrating practical robustness and extensible potential as parsing technology advances.
@@ 5 Conclusion
In this paper, we presented DISRetrieval, a discourse-aware hierarchical retrieval framework that systematically incorporates rhetorical structure theory into long document question answering. Our approach features three key innovations: language-universal discourse parsing, LLM-enhanced node representations, and structure-aware evidence selection. Comprehensive experiments on QASPER, QuALITY, NarrativeQA, and MultiFieldQA-zh demonstrate substantial improvements over existing methods across varying context lengths, embedding models, generation architectures, and languages. Ablation studies confirm that discourse structure provides more principled document organization than semantic clustering approaches, with the framework naturally adapting to different document types and information needs through multi-granularity retrieval. Our cross-lingual experiments validate that linguistically-grounded discourse structures effectively transcend language boundaries, demonstrating robust generalization from English to Chinese documents. Beyond performance improvements, this work demonstrates that linguistically-grounded approaches remain valuable in the era of LLMs, offering principled alternatives to purely data-driven methods. Our framework opens new directions for incorporating structured linguistic knowledge into neural retrieval systems, with potential applications in legal document analysis, scientific literature review, and educational content processing.
@@
While DISRetrieval demonstrates substantial improvements across multiple datasets and architectures, we acknowledge several limitations that point to promising future directions. First, our discourse parser’s performance inherently bounds the overall system effectiveness, as shown in Figure 4, where parser quality directly impacts downstream performance. However, our language-universal parser, trained with LLM-based multilingual data augmentation, demonstrates effective cross-lingual generalization from English to Chinese and consistent cross-genre improvements (research papers, fiction, movie scripts), showing practical robustness and extensible potential as parsing technology advances. While we currently demonstrate effectiveness on English and Chinese, extending to more languages would require additional translated training data through similar augmentation strategies. We note that the scarcity of suitable long-document QA datasets in other languages currently limits broader multilingual evaluation. Future work will explore expanding language coverage through collaboration with multilingual NLP communities, developing domain-adaptive parsers, and investigating multi-granularity discourse analysis. Second, our adaptive summarization strategy using threshold $\tau$ is relatively simple but proves effective across diverse document types ($\tau$=0 for academic papers, $\tau$=50 for narratives). While this approach successfully balances computational efficiency with representation quality, more sophisticated dynamic thresholding based on content complexity and hierarchical position could further optimize the trade-off. Third, current evaluation metrics, though comprehensive across four challenging datasets spanning multiple languages, may not fully capture the nuanced benefits of discourse-aware retrieval, such as structural coherence preservation and hierarchical information flow. We plan to develop evaluation frameworks that better assess discourse-aware retrieval quality across multiple dimensions and languages.
Ethical Considerations
@@
This research adheres to ethical principles in natural language processing research and does not raise significant ethical concerns. Our work focuses on improving document retrieval and question answering systems through discourse-aware hierarchical modeling, which has potential positive societal impacts by enhancing information access and comprehension across multiple languages. The datasets used in our experiments (QASPER, QuALITY, NarrativeQA, and MultiFieldQA-zh) are publicly available benchmark datasets that have been previously vetted by the research community and do not contain sensitive personal information or harmful content. Our discourse parsing approach operates on linguistic structures rather than content semantics, minimizing risks of bias amplification or misuse. The language-universal parser is trained using LLM-based translation for data augmentation, which may inherit potential biases from the translation model, though our focus on structural discourse relations rather than semantic content helps mitigate such risks. The proposed DISRetrieval framework is designed as a general-purpose retrieval enhancement technique that can benefit various applications requiring long document understanding across languages, such as academic research assistance, educational content processing, and legal document analysis. We acknowledge that like any information retrieval system, our method could potentially be misused if applied to spread misinformation or manipulate access to information, but such concerns are not specific to our approach and apply broadly to information retrieval technologies. We encourage responsible deployment of our system with appropriate safeguards and human oversight, particularly in high-stakes applications and cross-lingual scenarios. All experiments were conducted using publicly available computational resources and open-source tools, ensuring reproducibility and transparency in our research process.
@@ B.1 Dataset Specifications
MultiFieldQA-zh test 200 200 6701 14918
Table 7: Detailed information of the datasets used in our experiments.
@@
The NarrativeQA dataset is a significant advancement in evaluating machine reading comprehension. Unlike traditional datasets that focus on short passages, NarrativeQA emphasizes understanding long-form narratives, such as entire books or full-length movie scripts. One of the defining features of NarrativeQA is the length of its documents. On average, each document contains 51,372 words. Notably, some individual documents extend up to 346,902 words, making it one of the longest reading comprehension datasets available.
The MultiFieldQA-zh Chinese dataset comprises long-form documents from various sources, including legal texts, government reports, encyclopedias, and academic papers, each accompanied by relevant questions and answers. Compared to the English documents, the Chinese data exhibits unique linguistic structures, expressions, and domain-specific terminology, posing higher demands on models’ capabilities for long-text comprehension and cross-domain information integration. In our experiments, we use the test set for evaluation and comparison. More detailed information is presented in Table 7.
B.2 RST-DT Data Translation
To train a cross-lingual discorse parser, we translate the RST-DT training data into Chinese to augment the available Chinese data. Given the syntactic and word-order differences across languages, we perform translation at the sentence level, which is also consistent with our training objective of sentence-level RST parsing. We then combine the original RST-DT data, the sentence-level RST data, and the Chinese-translated data to form the final training corpus. For the translation process, we employ the GPT-4o-mini model, and the prompt is provided below:
Prompt for Translation
System: You are a helpful assistant.
User: Please translate the following sentences into Chinese. Translate each sentence individually while preserving the original order, and return the results in JSON format.
For example:
Original sentences:
1. Sentence 1
2. Sentence 2
3. Sentence 3
Translation:
{1: translated sentence, 2: translated sentence, …}
Strictly return the output in valid JSON format. Do not include any additional text, and ensure that the output can be directly parsed as JSON.
Original sentences: {sentences}
LLM: JSON-formatted Chinese translations.
B.3 Model Specifications
@@ B.4 Implementation Details
B.5 Experiments Compute Resources
@@ B.6 Prompts
Given text: {left child node text} {input child node text}
@@
User: Using the following information: {context}. Answer the following question in less than 5-7 words, if possible. For yes or no question, only return ’yes’ or ’no’.
@@ C.1 Computational Efficiency and Scalability
Method 10K words 30K words 50K words 70K words 90K words
RAPTOR 63.4s 194.3s 338.5s 488.0s 611.5s
DISRetrieval 20.3s 68.1s 103.0s 124.4s 159.9s
[…]
Table 8: Computational efficiency comparison between DISRetrieval and RAPTOR using NVIDIA A800 GPU and LLaMA3.1-8B-Instruct.
To provide a comprehensive evaluation of computational efficiency, we conducted a direct comparison between our method and RAPTOR, another preprocessing-based hierarchical retrieval approach. Both methods were evaluated under identical hardware and software configurations to ensure fair comparison. Specifically, we used a single NVIDIA A800 GPU with LLaMA3.1-8B-Instruct as the language model for both discourse tree construction in our method and recursive summarization in RAPTOR.
[…]
The computational efficiency of our method stems from two key factors. First, discourse parsing operates at the sentence level rather than requiring recursive processing of all text segments, which significantly reduces the number of LLM calls needed. Second, our LLM enhancement process is selective and targeted, focusing only on internal nodes that require summarization, whereas RAPTOR performs recursive summarization across all levels of the hierarchy. Despite this substantial speedup, our method achieves superior retrieval performance as demonstrated in Tables 1 2 4, indicating that computational efficiency does not come at the cost of effectiveness.
It is important to note that both discourse parsing and LLM enhancement in our framework are query-independent preprocessing steps performed once per document. In real-world deployment scenarios where multiple queries are issued against the same document collection, these preprocessing costs are amortized across all queries, making the per-query computational overhead negligible. This characteristic makes our approach particularly suitable for applications such as scientific literature review, legal document analysis, and enterprise knowledge management, where document collections are relatively stable but query patterns are diverse and frequent.
@@ C.4 Ablation study on parameter $\tau$ across different datasets
Input: Document $D$ containing sentences $S_{i}$, paragraphs $P_{i}$; Query $q$
Output: Retrieved evidence segments $E$
/* Stage 1: Discourse-Aware Tree Construction */
for each paragraph $P_{i}\in D$ do
   Initialize stack $\sigma$ and sentence queue $\beta\leftarrow S_{i}$;
 while $\beta$ not empty OR $|\sigma|>1$ do
    /* RST parsing operations */
[…]
    // Determine next parsing action
    switch $a_{t}$ do
       case shift do
            Move next sentence from $\beta$ to $\sigma$;
          // Add new sentence to stack
         end case
       case reduce do
          $s_{2},s_{1}\leftarrow$ Pop top two elements from $\sigma$;
            Determine discourse relation $r$ between $s_{1}$ and $s_{2}$;
          // Using RST relations
            Create new node with relation $r$;
            Push new node to $\sigma$;
         end case
       case pop_root do
            Final tree node $\leftarrow$ Pop from $\sigma$;
          // Complete the tree
         end case
      end switch
   end while
 $T_{i}\leftarrow$ Resulting discourse tree;
 /* LLM Enhancement for Paragraph Tree */
 for each non-leaf node $v\in T_{i}$ (bottom-up order) do
    $v_{l},v_{r}\leftarrow$ Get left and right children;
    if $|v_{l}|+|v_{r}|\geq\tau$ then
       $v^{*}\leftarrow f_{\text{LLM}}(v_{l},v_{r})$;
       // Generate concise summary using LLM
    else
       $v^{*}\leftarrow f_{merge}(v)$;
       // Merge the subtree rooted at v
      end if
   end for
  Store root text representation $t^{i}_{\text{root}}$ for $T_{i}$;
end for
Initialize stack $\sigma$ and queue $\beta\leftarrow\{t^{1}_{\text{root}},...,t^{n}_{\text{root}}\}$;
while $\beta$ not empty OR $|\sigma|>1$ do
 /* Similar RST parsing at document level */
   Apply RST parsing operations as in Phase 1;
   Build document tree $T_{\text{doc}}$;
end while
/* Stage 2: Node Representation */
for each non-leaf node $v\in T_{\text{doc}}$ (bottom-up order) do
 $t_{l},t_{r}\leftarrow$ Get content from left and right children;
 if $|t_{l}|+|t_{r}|\geq\tau$ then
    $t_{v}\leftarrow f_{\text{LLM}}(t_{l},t_{r})$;
    // LLM-based enhancement for document-level nodes
 else
    $t_{v}\leftarrow t_{l}\oplus t_{r}$;
    // Direct concatenation for short text, don’t need to merge subtree
   end if
end for
$T_{D}\leftarrow$ Replace leaf nodes in $T_{\text{doc}}$ with corresponding $T_{i}$;
// Integrate trees into unified structure
for each node $v\in T_{D}$ do
 $e_{v}\leftarrow\text{Encoder}(t_{v})$;
 // Generate dense vector representations
end for
/* Stage 3: Hierarchical Evidence Retrieval and Selection */
$e_{q}\leftarrow\text{Encoder}(q)$;
// Transform query to embedding space
$\text{scores}\leftarrow\{\}$, $E\leftarrow\{\}$;
for node $v\in T_{D}$ do
 $\text{scores}[v]\leftarrow\text{cosine}(e_{q},e_{v})$;
 // Compute relevance scores
end for
Apply Algorithm 1 for evidence selection;
return $E$
Algorithm 2 DISRetrieval: Discourse Structure-based Long Document Retrieval